Find the angles of triangles in which a=2,b=1and c=√3

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
Cos A = b² + c² - a² /2bc = 4-4 /2bc = 0
 
Hence
 
Angle A = 90°
 
Cos B = c² + a² - b² /2ac = 6 /4 sqrt(3) = (sqrt 3)/2
 
Hence B = 30 °
 
Now C = 60°
 
 
Regards
Arun